

********************************************************************
UARS WEEKLY STATUS REPORT - 13 APR 2001
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, taking solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. We continue to gather enough data to meet all the SOLSTICE
science objectives during this period of reduced tape recorder operations.
***********
SUSIM
***********
Nothing received.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
PEM/AXIS was activated in its full power mode on April 1 and deactivated
on April 9. At the current time, AXIS is not generating energy deposition.
SwRI is looking into this. Tests of AXIS production code on the
rehost computer are continuing.
At the present time, PEM energy deposition from its spectrometers is not
being generated. Recent data configuration changes have affected PEM software
causing multiple failures. These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update
processing code with a patch delivery. The delivery patch may include changes
which would allow the PEM software to execute on a DEC Alpha machine. GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha).
PEM has examined its processed files of the rehost machine and has found
that no differences exist in the Level 1 and Level 2 output files which were
generated the rehost when the same input data was used on CDHF and CDHFB.
PEM is now confident that output generated by PEM 3.32, Jobs 1 & 2, on the
rehost system is valid and correct.
The PEM team is working on replacing the PEM:: machine with a DEC Alpha
machine to be called PEM2K::. Tests of the NRT data transfer from CDHF2::
through PEM2K:: to the unix computer running the PEM GSE software show no
difference between equivalent NRT files transferred from PEM::. Thus, NRT
data from CDHF2:: is captured correctly on PEM2K::. Tests of the NRT data
transfer from CDHFB:: were not successful from both PEM:: and PEM2K::.
Results of the test of NRT data transfers from the rehost computer were
successful for individual files which exist on CDHFB::. PEM will participate
in the NRT data flow to capture NRT data from CDHFB:: when specified by GSFC.
PEM has been working with GSFC on allowing commanding of the PEM instrument
by creating DAPs on the PEM2K machine and submitting them to the CMS.
So far, this has not been achieved. It was discovered that PEM2K is
running Openwindows version 7.2-1 which operates differently across the
network than Openwindows version 7.1-2 (the CDHFB operating system). Compaq
is investigating to determine a solution. In the mean time, Steve Bove made
some suggestions which seem to have improved the PEM2K:: to UCMS:: image
execution ability; however, the CMS has had troubles recently and it appears
that the UCMS:: definitions for PEM2K:: must not be correct on the UCMS:: node.
PEM is waiting on a reply back from GSFC.
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
HRDI continues to function normally, making wind measurements during
the sunlit parts of orbits nearly every day, and measuring nighttime
winds when power is available.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Following repair of the UARS Command Management System
on 2 April, daily processing of instrument Daily Activity
Plans continues with no problems. Instrument operations
schedule is now complete to 30 May 2001.
***********
CDHF
***********
CURRENT SYSTEM
---------------
The CDHF and CDHF REHOST will be unavailable from April 22 9:30 GMT until
power up time at 11:30 GMT (tentative) April 23. This is due to a scheduled
building power outage.
This information is contain in the Login Notice as well as a telephone outage
schedule.
UARS CDHF REHOST
-----------------
Validating a mandatory upgrade of the HSM. This is the software that "Puts"
and Gets" data from the Silo.
The current system is still the "real" system until remaining issues are
resolved. This should be about another week. Production processing results are
not changed with this upgrade.
***********
FOT
***********
REPORTING PERIOD
This report covers 31 March 2001 (Orbit 52216, GMT Day 090) through 06 April
2001 (Orbit 52319, GMT Day 096).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 090/1931 52227 Power OFF
HALOE 090/2111 52228 Power ON
PEM/AXIS 091/0231 52232 Power ON
HALOE 092/0104 52246 Power OFF (HI TBAT)
HALOE 092/0232 52247 Power ON
HALOE 093/0100 52261 Power OFF (HI TBAT)
HALOE 093/0233 52262 Power ON
HALOE 093/2333 52275 Power OFF
HALOE 094/0118 52276 Power ON
HALOE 095/0415 52293 Power OFF
HALOE 095/0554 52294 Power ON
The beta angle ranged between the angles of 50.2 degrees to 48.1 degrees for
this report period. The beta angle is now increasing towards a maximum angle
of 74 degrees on 09 May 2001 (DOY 129).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.3 and 1.6 degrees C. Additional battery performance data for each
battery on 30 March 2001 and 06 April 2001 are provided in Appendix D.
The clock error ranged from Ð10.9 to 10.4 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 00
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 25
¥ PEM/AXIS = 02
¥ HALOE = 10
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 03
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/316 - 2287.496493
· Transponder B: 01/087 - 2287.499152
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 01/067 show the observatory data loss to be 300 hours, 1
minute, 12.760 seconds. This shows no change since the last report period. This
is a 0.39 percent data loss, which equals a 99.61 percent data capture for the
mission.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
UARS is using both the Fine Sun Sensor (FSS) and the Three Axis Magnetometer
(TAM) as inputs to UFLTR this week. The FOT is pleased with their performance,
but continues to make adjustments to the TAM alignment to improve its
performance and UARS's attitude.
The FOT awaits the completion of the FSW patch that will improve the magnetic
field model from a 5th order to 7th order solution. The patch is currently
beeing tested on the UARS Testing and Training Simulator (UTTS). Testing should
be completed in the next few weeks.
MEETINGS
The FOT attended the UARS Monthly Meeting on Monday, April 04, 2001.
The FOT attended a UARS Flight Software Meeting on Thursday, March 05, 2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 19 April 2001 (DOY 109). This maneuver
will be a Reverse to Forward Flight maneuver (Southern Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 06 April 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
52217 090/0229 01-193 NBTR NBTR PBK/RTS-58
52220 090/0729 01-193 NBTR NOOPS D-090
52227 090/1834 01-195 NBTR NOOPS D-091
52227 090/1931 01-196 HALOE Power OFF
52228 090/2111 01-197 HALOE Power ON
52232 091/0231 01-195 NBTR NBTR PBK/RTS-58
52232 091/0231 01-198 PEM/AXIS Power ON (BEGIN)
52233 091/0412 01-198 PEM/AXIS Power ON (COMPLETE)
52235 091/0729 01-195 NBTR NOOPS D-091
52242 091/1835 01-199 NBTR NOOPS D-092
52244 091/2114 01-200 OBC Clk Adj 20.6 to 0.00 msec/day
52246 092/0104 01-201 HALOE Power OFF (HI TBAT)
52247 092/0232 01-202 HALOE Power ON
52247 092/0232 01-199 NBTR NBTR PBK/RTS-58
52250 092/0729 01-199 NBTR NOOPS D-092
52257 092/1836 01-203 NBTR NOOPS D-093
52261 093/0100 01-204 HALOE Power OFF (HI TBAT)
52262 093/0233 01-203 NBTR NBTR PBK/RTS-58
52262 093/0233 01-205 HALOE Power ON
52265 093/0730 01-203 NBTR NOOPS D-093
52274 093/2152 01-206 OBC/CLK Clk Adj 20.6 to 0.00 msec/day
52275 093/2333 01-207 HALOE Power OFF
52276 094/0118 01-208 HALOE Power ON
52277 094/0235 01-209 NBTR NBTR PBK/REC
52280 094/0817 01-209 NBTR NOOPS D-094
52281 094/0956 01-209 NBTR NBTR PBK/RTS-58
52284 094/1432 01-210 MACS TAM CALIBRATION
52286 094/1755 01-211 NBTR NOOPS D-095
52292 095/0236 01-211 NBTR NBTR PBK/RTS-58
52293 095/0415 01-212 HALOE Power OFF
52294 095/0554 01-213 HALOE Power ON
52295 095/0732 01-211 OBC/NBTR NOOPS D-095
52301 095/1755 01-214 MACS TAM CAL.
52301 095/1755 01-215 NBTR NOOPS D-096
52308 096/0416 01-216 OBC Clk Adj fr 0.00 to 20.6 msec/day
52309 096/0636 01-215 NBTR Day 096 noops
52316 096/1756 01-217 NBTR Day 097 noops
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 06 April 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 06 April 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
52244 091/2117 Clk Adj from 20.6 to 0.00 msec/day
52274 093/2152 Clk Adj from 20.6 to 0.00 msec/day
52308 096/0419 Clk Adj from 0.00 to 20.6 msec/day
___________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 06 April 2001
30 Mar 2001 - GMT Day 089 Beta = 48.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- ------- ---------- ----- ------- ------
1 +039.2/+011.2 -3.05/-4.62 +0.0/+0.0 13.0 0.000 0.0
2 +22.4/-67.2 +5.15/+3.87 +26.4/-15.2 25.0 1.050 19.6
3 +22.4/-44.8 +4.51/+2.92 +25.6/-15.2 25.0 1.050 19.3
06 Apr 2001 - GMT Day 096 Beta = 48.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- ------- ---------- ----- ------- ------
1 +044.8/+022.4 -3.68/-4.62 +0.0/+0.0 13.0 0.000 0.0
2 +22.4/-44.8 +5.15/+3.87 +26.4/-15.2 25.0 1.050 19.7
3 +22.4/-33.6 +3.87/+2.60 +25.6/-15.2 25.0 1.040 19.5
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
___________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 06 April 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-014 FOT 52225 090/1608 14170 Generic Late Acquisition
01-019 FOT 52315 096/1616 14170 Generic Late Acquisition
01-020 FOT 53318 096/1946 23815 Handover CSC
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-014 FOT see list Mar 01 14170 Generic Late Acquisition
01-020 FOT 53318 096/1946 23815 Handover CSC
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-019 FOT see list Apr 01 14170 Generic Late Acquisition
SEVNTFW01/14